OpenStudy (anonymous):

solve the system of linear equations by graphing. [3x+y=0 [2x+y=2

OpenStudy (anonymous):

get y by its self do u know how to do that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

okay well you divide 3x by 0 and what do u get?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

no for equations like this u just keep it in the equation so what you suppose to get is y=3x+0 and now you have something to graph

OpenStudy (anonymous):

no try and do it on the next one

OpenStudy (anonymous):

i meant now try*

OpenStudy (anonymous):

so is it -3x ?

OpenStudy (anonymous):

no where did u get the negative from?

OpenStudy (anonymous):

i gave u the answer now all u have to do it graph it

OpenStudy (anonymous):

but u still have to get the 2x+y=2 try to do it

OpenStudy (anonymous):

I don't know how to graph that

OpenStudy (anonymous):

the answer is [y=3x+0 [y=2x+2 you dont know how to graph that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

do u know what the y-intercept is?

OpenStudy (anonymous):

and the slope?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

do you know this equation y=mx+b?

OpenStudy (anonymous):

m is the slope and b is the y-intercept?

OpenStudy (anonymous):

I don't understand :(

OpenStudy (anonymous):

is this your homework?

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

u have \[ \large 3x+y=0\qquad\text{and}\qquad 2x+y=2 \]

OpenStudy (anonymous):

ok

OpenStudy (helder_edwin):

isolating y u have \[ \large y=-3x\qquad\text{and}\qquad y=-2x+2 \] ok?

OpenStudy (anonymous):

ok

OpenStudy (helder_edwin):

the easiest way to graph lines is doing a table. u need only two points.

OpenStudy (anonymous):

ok

OpenStudy (helder_edwin):